OpenStudy (anonymous):

Problem: Two cars A and B are 200km apart and they are traveling toward each other at the speeds of 28km/h and 22 km/h respectively. When and where will they meet? 2. If the car A and B in problem 1 start from the same point, going toward the same direction, at what time will they be 18km apart? 3. If B starts at 1:00pm and gives chase starting at 1:15 from the same point as B, when and where will A overtake B? Thank you so much for answering!!

OpenStudy (anonymous):

Help...

OpenStudy (anonymous):

1) relative speed =28+22=50 they will meet after 200/50=4 hr at a distance of 28*4=112 km from the starting poing or (200-112=88 from the other sid)

OpenStudy (anonymous):

2) relative speed now =28-22=6 they will be 18 km apart after 18/6 = 3 hr

OpenStudy (anonymous):

@huneya : you just copied matricked's answers..... now see, that is my problem, i don't get how matricked did it....

OpenStudy (anonymous):

i m not copying anyone thats the answers

OpenStudy (anonymous):

like how did you get 18???

OpenStudy (anonymous):

in ur second qusetion its says 18km

OpenStudy (anonymous):

hey alynamarie in 3rd question does A also starts at same point A ?

OpenStudy (anonymous):

@alynamarie read ur question before

OpenStudy (anonymous):

@clark_boy : i dunno, that's why i need help understanding this stuff

OpenStudy (anonymous):

@huneya : yup, i read it already, but still can's understand

OpenStudy (phi):

Here is one way to think of part 1. remember distance= rate * time